OpenStudy (anonymous):

3(5-2x)=-2(6-3x)-10x help please

OpenStudy (academicgurusinc):

3(5-2x)=-2(6-3x)-10x 15 -6x =-12 + 6x -10x 27= 2x x= 27/2 We hope this helps. Regards, Academic GurusInc. Twitter (@Academic_Gurus) Facebook (AcademicGurusInc)